1 point 5л Evaluate csc without using a calculator. Be 6 sure to show scratch work. -0.5 2 OA OD 0.5 -2 С OB Q
Answer #1

CSC 5 6 = CSC (65-T (6***) =csc ( 67 - 7 CSC(T-7 = CSc(2x7 - 7 Sin All + tan COS - & CSCA 1 [sin 7 = 1 because Sint 6 rs 2 2